[31 Cash Conversion Cycle (20 points) (a) Calculate the Cash Conversion Cycle (CCC) for Porto Electronics in 2018 with the followin information. Note: use the information from receivables, inventory and payables at start. Please show yo workings. From the balance sheet Receivables 2018 . Receivables 2017 Inventory 2018 Inventory 2017 30,000 15,000 32,540 28,560 27,000 22,000 . Payables 2018 Payables 2017 From the 2018 income statement Sales 430,700 COGS 328,500 (b) What would be the effect on the CCC of an improvement in the efficiency of Porto El inventory management. Would it increase? Would it decrease? Briefly explain why.
Answer #1

HI,

The formula for Cash conversion cycle = Days inventory outstanding + Days Sale Outstanding -Days payable outstanding

Here Days Inventory outstanding (DIO)= Average inventory/COGS per day

= (32540+28560)/2/328500/365

= 30550*365/328500 = 33.94 days

Same way Days Sales Outstanding DSO = Average Receivable/Sales per day

= (30000+15000)/2/430700/365 = 19.07 days

Days payable outstanding (DPO) = Average Payable/COGS per day

(27000+22000)/2/328500/365 = 27.22 days

So Cash Conversion cycle CCC = 33.94 + 19.07 - 27.22 = 25.79 days

That means firm takes 25.79 to generate its cash from working capital.

b) If the inventory management will be more efficient that means firm's will maintain lesser inventory in its hand and so its Days inventory outstanding will reduce and with that its CCC also will come down,

Thanks